OpenStudy (anonymous):

2. Which of the following would be a reasonable estimate for the height of a skyscraper? a 1 × 10squared by -8 feet b 1 × 10squared by 8 feet c 1 × 10squared by -3 feet d 1 × 10squared by 3 feet

OpenStudy (anonymous):

B

OpenStudy (anonymous):

and it's not "squared by", it's "to the power of"

OpenStudy (tkhunny):

It is not "squared by". "squared" applies ONLY to the number 2. You can notate exponentiation with this guy "^". a 1 × 10^-8 feet b 1 × 10^8 feet c 1 × 10^-3 feet d 1 × 10^3 feet That being said, give it a little thought with the understanding of the notation. a 1 × 10^-8 feet = 0.00000001 ft = 0.00000012 in = 30.48 Angstroms - VERY SHORT building. Actually, rather large for an atom. b 1 × 10^8 feet = 100000000 ft = 18,939 miles - It will not get you to the moon, but you ARE on your way. c 1 × 10^-3 feet = 0.001 ft = 0.012 in - Still pretty short, but you can SEE this one. d 1 × 10^3 feet = 1000 ft - That's about 2/3 of the Empire State Building. I'm going with D.